Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Data collection)
letendre, masserot - 16:58 Wednesday 15 March 2023 (59265) Print this report
WI Suspension timing issues

Today each CEB suspension crate has its owm Timing Distribution box (TDB) receiving by fibers the IRIGB signal and 10MHz clock from the DAQ room , then inside a suspension rack  the Timing signals are propaged using RJ45 cables to the DSP boards .

At the DAQ  room  level (scheme),

  • the TDB05B  provides the timing optical signals to the NI, SR, BS and BS suspensions
  • the TDB06B provides the timing optical signals to the  WI, IB and OB suspensions, to the DBOX located in the EEROM

Today there is not a direct monitoring  of the Timing loop stability of the suspension DSPs . An indirect monitoring is performed using the emitted  GPS time stamp of the Tolm packet  with the channels Daq_S{a,c}_XX_Tolm_Send_Start_min.

The attached plots shows the differents Daq_S{a,c}_XX_Tolm_Send_Start_min signals,  the timing _error signals of the DAQ boxes located in the EEROOM, the DBOX located in the CEB platform  and the monitoring of the IRIGB pulse at the level of the first TDB04B .  The red rectangle highlight the suspensions where there was some  timing jumps : Sc_BS, S{a,c}_NI, and S{a,c}_WI. There was 2 jumps space by 4mn . The jump level was not the same on all these suspensions, the big jumps occurred only on the WI suspension DSPs .

it seems unlikely that the troubles occurred in the fiber propagation , it 's not obvious to find where these glitches comes from but the only way to recover the correct situtation it is to reconfigure the Suspension timing part., at least the WI one

Images attached to this report
Non-image files attached to this report
Comments to this report:
Boschi - 17:06 Wednesday 15 March 2023 (59267) Print this report

Both Sa and Sc WI master boards have been reset. Timing synchronization seems correct now.

masserot - 18:19 Wednesday 15 March 2023 (59268) Print this report

The plots shows the prevoius channels after the recovery : the Daq_S{a,c}_WI_Tolm_Send_Start_min channels returns to theirs previous values, as well the Daq_Sc_NI_Tolm_Send_Start_min one .

Only the WI DSPs were resynchronized but it seems that this has an impact on the Sc_NI DSP cycles too

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×